$SPAD/src/input gary1.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(curve(cos(4*t)*cos(7*t),cos(4*t)*sin(7*t)),t = 0..2*%pi,[title "Greetings!"])

———-

References^{[1] nothing}

4